\begin{thebibliography}{26}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Abernethy et~al.(2008)Abernethy, Bartlett, Rakhlin, and
  Tewari]{Minimax:Online}
J.~Abernethy, P.~L. Bartlett, A.~Rakhlin, and A.~Tewari.
\newblock Optimal stragies and minimax lower bounds for online convex games.
\newblock In \emph{Proceedings of the 21st Annual Conference on Learning
  Theory}, pages 415--423, 2008.

\bibitem[Besbes et~al.(2015)Besbes, Gur, and Zeevi]{Non-Stationary}
O.~Besbes, Y.~Gur, and A.~Zeevi.
\newblock Non-stationary stochastic optimization.
\newblock \emph{Operations Research}, 63\penalty0 (5):\penalty0 1227--1244,
  2015.

\bibitem[Boyd and Vandenberghe(2004)]{Convex-Optimization}
S.~Boyd and L.~Vandenberghe.
\newblock \emph{Convex Optimization}.
\newblock Cambridge University Press, 2004.

\bibitem[Cesa-Bianchi and Lugosi(2006)]{bianchi-2006-prediction}
N.~Cesa-Bianchi and G.~Lugosi.
\newblock \emph{Prediction, Learning, and Games}.
\newblock Cambridge University Press, 2006.

\bibitem[Chiang et~al.(2012)Chiang, Yang, Lee, Mahdavi, Lu, Jin, and
  Zhu]{Gradual:COLT:12}
C.-K. Chiang, T.~Yang, C.-J. Lee, M.~Mahdavi, C.-J. Lu, R.~Jin, and S.~Zhu.
\newblock Online optimization with gradual variations.
\newblock In \emph{Proceedings of the 25th Annual Conference on Learning
  Theory}, 2012.

\bibitem[Daniely et~al.(2015)Daniely, Gonen, and
  Shalev-Shwartz]{Adaptive:ICML:15}
A.~Daniely, A.~Gonen, and S.~Shalev-Shwartz.
\newblock Strongly adaptive online learning.
\newblock In \emph{Proceedings of the 32nd International Conference on Machine
  Learning}, pages 1405--1411, 2015.

\bibitem[Duchi et~al.(2011)Duchi, Hazan, and Singer]{JMLR:Adaptive}
J.~Duchi, E.~Hazan, and Y.~Singer.
\newblock Adaptive subgradient methods for online learning and stochastic
  optimization.
\newblock \emph{Journal of Machine Learning Research}, 12:\penalty0 2121--2159,
  2011.

\bibitem[Hall and Willett(2013)]{Dynamic:ICML:13}
E.~C. Hall and R.~M. Willett.
\newblock Dynamical models and tracking regret in online convex programming.
\newblock In \emph{Proceedings of the 30th International Conference on Machine
  Learning}, pages 579--587, 2013.

\bibitem[Hazan(2016)]{Intro:Online:Convex}
E.~Hazan.
\newblock Introduction to online convex optimization.
\newblock \emph{Foundations and Trends in Optimization}, 2\penalty0
  (3-4):\penalty0 157--325, 2016.

\bibitem[Hazan and Seshadhri(2007)]{Adaptive:Hazan}
E.~Hazan and C.~Seshadhri.
\newblock Adaptive algorithms for online decision problems.
\newblock \emph{Electronic Colloquium on Computational Complexity}, 88, 2007.

\bibitem[Hazan et~al.(2007)Hazan, Agarwal, and Kale]{ML:Hazan:2007}
E.~Hazan, A.~Agarwal, and S.~Kale.
\newblock Logarithmic regret algorithms for online convex optimization.
\newblock \emph{Machine Learning}, 69\penalty0 (2-3):\penalty0 169--192, 2007.

\bibitem[Herbster and Warmuth(1998)]{Herbster1998}
M.~Herbster and M.~K. Warmuth.
\newblock Tracking the best expert.
\newblock \emph{Machine Learning}, 32\penalty0 (2):\penalty0 151--178, 1998.

\bibitem[Hoeffding(1963)]{Hoeffding}
W.~Hoeffding.
\newblock Probability inequalities for sums of bounded random variables.
\newblock \emph{Journal of the American Statistical Association}, 58\penalty0
  (301):\penalty0 13--30, 1963.

\bibitem[Jadbabaie et~al.(2015)Jadbabaie, Rakhlin, Shahrampour, and
  Sridharan]{Dynamic:AISTATS:15}
A.~Jadbabaie, A.~Rakhlin, S.~Shahrampour, and K.~Sridharan.
\newblock Online optimization: Competing with dynamic comparators.
\newblock In \emph{Proceedings of the 18th International Conference on
  Artificial Intelligence and Statistics}, pages 398--406, 2015.

\bibitem[Jun et~al.(2017)Jun, Orabona, Wright, and
  Willett]{Improved:Strongly:Adaptive}
K.-S. Jun, F.~Orabona, S.~Wright, and R.~Willett.
\newblock Improved strongly adaptive online learning using coin betting.
\newblock In \emph{Proceedings of the 20th International Conference on
  Artificial Intelligence and Statistics}, pages 943--951, 2017.

\bibitem[Mokhtari et~al.(2016)Mokhtari, Shahrampour, Jadbabaie, and
  Ribeiro]{Dynamic:Strongly}
A.~Mokhtari, S.~Shahrampour, A.~Jadbabaie, and A.~Ribeiro.
\newblock Online optimization in dynamic environments: Improved regret rates
  for strongly convex problems.
\newblock In \emph{IEEE 55th Conference on Decision and Control}, pages
  7195--7201, 2016.

\bibitem[Rakhlin and Sridharan(2013)]{Predictable:COLT:2013}
A.~Rakhlin and K.~Sridharan.
\newblock Online learning with predictable sequences.
\newblock In \emph{Proceedings of the 26th Conference on Learning Theory},
  pages 993--1019, 2013.

\bibitem[Shalev-Shwartz(2011)]{Online:suvery}
S.~Shalev-Shwartz.
\newblock Online learning and online convex optimization.
\newblock \emph{Foundations and Trends in Machine Learning}, 4\penalty0
  (2):\penalty0 107--194, 2011.

\bibitem[Shalev-Shwartz et~al.(2007)Shalev-Shwartz, Singer, and
  Srebro]{ICML_Pegasos}
S.~Shalev-Shwartz, Y.~Singer, and N.~Srebro.
\newblock Pegasos: primal estimated sub-gradient solver for {SVM}.
\newblock In \emph{Proceedings of the 24th International Conference on Machine
  Learning}, pages 807--814, 2007.

\bibitem[Srebro et~al.(2010)Srebro, Sridharan, and Tewari]{NIPS2010_Smooth}
N.~Srebro, K.~Sridharan, and A.~Tewari.
\newblock Smoothness, low noise and fast rates.
\newblock In \emph{Advances in Neural Information Processing Systems 23}, pages
  2199--2207, 2010.

\bibitem[van Erven and Koolen(2016)]{NIPS2016_6268}
T.~van Erven and W.~M. Koolen.
\newblock Metagrad: Multiple learning rates in online learning.
\newblock In \emph{Advances in Neural Information Processing Systems 29}, pages
  3666--3674, 2016.

\bibitem[Yang et~al.(2016)Yang, Zhang, Jin, and Yi]{Dynamic:2016}
T.~Yang, L.~Zhang, R.~Jin, and J.~Yi.
\newblock Tracking slowly moving clairvoyant: Optimal dynamic regret of online
  learning with true and noisy gradient.
\newblock In \emph{Proceedings of the 33rd International Conference on Machine
  Learning}, pages 449--457, 2016.

\bibitem[Zhang et~al.(2013)Zhang, Yi, Jin, Lin, and He]{ICML:13:Zhang:Sparse}
L.~Zhang, J.~Yi, R.~Jin, M.~Lin, and X.~He.
\newblock Online kernel learning with a near optimal sparsity bound.
\newblock In \emph{Proceedings of the 30th International Conference on Machine
  Learning}, 2013.

\bibitem[Zhang et~al.(2017)Zhang, Yang, Yi, Jin, and
  Zhou]{Dynamic:Regret:Squared}
L.~Zhang, T.~Yang, J.~Yi, R.~Jin, and Z.-H. Zhou.
\newblock Improved dynamic regret for non-degenerate functions.
\newblock In \emph{Advances in Neural Information Processing Systems 30}, pages
  732--741, 2017.

\bibitem[Zhang et~al.(2018)Zhang, Yang, Jin, and Zhou]{Dynamic:Regret:Adaptive}
L.~Zhang, T.~Yang, R.~Jin, and Z.-H. Zhou.
\newblock Dynamic regret of strongly adaptive methods.
\newblock In \emph{Proceedings of the 35th International Conference on Machine
  Learning}, 2018.

\bibitem[Zinkevich(2003)]{zinkevich-2003-online}
M.~Zinkevich.
\newblock Online convex programming and generalized infinitesimal gradient
  ascent.
\newblock In \emph{Proceedings of the 20th International Conference on Machine
  Learning}, pages 928--936, 2003.

\end{thebibliography}

\documentclass{article}

 \usepackage[preprint]{nips_2018}

\usepackage[utf8]{inputenc} 
\usepackage[T1]{fontenc}    
\usepackage{hyperref}       
\usepackage{url}            
\usepackage{booktabs}       
\usepackage{amsmath,amsfonts}       
\usepackage{nicefrac}       
\usepackage{microtype}      

\usepackage{algorithm,algorithmic}

\usepackage{paralist}

\def \E {\mathrm{E}}
\def \x {\mathbf{x}}

\def \u {\mathbf{u}}
\def \H {\mathcal{H}}

\def \R {\mathbb{R}}

\def \F {\mathcal{F}}

\def \ub {\bar{\u}}

\def \Pb {\overline{P}}

\def \C {\mathcal{C}}

\def \m {\mathbf{m}}

\def \xb {\bar{\x}}

\def \F {\mathcal{F}}

\def \X {\mathcal{X}}

\DeclareMathOperator*{\argmin}{argmin}

\newtheorem{thm}{Theorem}

\newtheorem{lemma}{Lemma}

 \newtheorem{ass}{Assumption}

\title{Adaptive Online Learning in Dynamic Environments}

\author{
  Lijun Zhang, \quad Shiyin Lu, \quad Zhi-Hua Zhou \\
  National Key Laboratory for Novel Software Technology\\
  Nanjing University, Nanjing 210023, China \\
  \texttt{\{zhanglj, lusy, zhouzh\}@lamda.nju.edu.cn}
}

\begin{document}

 \maketitle

\begin{abstract}
In this paper, we study online convex optimization in dynamic environments, and aim to bound the dynamic regret with respect to \emph{any} sequence of comparators. Existing work have shown that online gradient descent enjoys an $O(\sqrt{T}(1+P_T))$ dynamic regret, where $T$ is the number of iterations and $P_T$ is the path-length of the comparator sequence.  However, this result is unsatisfactory, as there exists a large gap from the $\Omega(\sqrt{T(1+P_T)})$ lower bound established in our paper. To address this limitation, we develop a novel online method, namely adaptive learning for dynamic environment (Ader), which achieves an optimal $O(\sqrt{T(1+P_T)})$ dynamic regret. The basic idea is to maintain a set of experts, each attaining an optimal dynamic regret for a specific path-length, and combines them with an expert-tracking algorithm.  Furthermore, we propose an improved Ader based on the surrogate loss, and in this way the number of gradient evaluations per round is reduced from $O(\log T)$ to $1$. Finally, we extend Ader to the setting that a sequence of dynamical models is available to characterize the comparators.
\end{abstract}

\section{Introduction}
Online convex optimization (OCO) has become a popular learning framework for modeling various real-world problems, such as online routing, ad selection for search engines and spam filtering \citep{Intro:Online:Convex}.  The protocol of OCO is as follows: At iteration $t$, the online learner chooses  $\x_t$ from a convex set $\X$.  After the learner
has committed to this choice, a convex cost function $f_t:\X \mapsto \R$ is revealed. Then, the learner suffers an instantaneous loss $f_t(\x_t)$, and the goal is to minimize the cumulative loss over $T$ iterations. The standard performance measure of OCO is regret:
\begin{equation} \label{eqn:regret}
\sum_{t=1}^T f_t(\x_t) - \min_{\x \in \X} \sum_{t=1}^T f_t(\x)
\end{equation}
which is the cumulative loss of the learner minus that of the best constant point chosen in hindsight.

The notion of regret has been extensively studied, and there exist plenty of algorithms and theories for minimizing regret \citep{ICML_Pegasos,ML:Hazan:2007,NIPS2010_Smooth,JMLR:Adaptive,Online:suvery,ICML:13:Zhang:Sparse}. However, when the environment is changing, the traditional regret  is no longer a suitable measure, since it compares the learner against a \emph{static} point. To address this limitation, recent advances in online learning have introduced an enhanced  measure---\emph{dynamic} regret, which received considerable research interest over the years \citep{Dynamic:ICML:13,Dynamic:AISTATS:15,Dynamic:Strongly,Dynamic:2016,Dynamic:Regret:Squared}.

In the literature, there are two different forms of dynamic regret. The general one is introduced by \citet{zinkevich-2003-online}, who proposes to compare the cumulative loss of the learner against \emph{any} sequence of comparators
\begin{equation} \label{eqn:dynamic:1}
R(\u_1, \ldots, \u_T) =\sum_{t=1}^T f_t(\x_t)  - \sum_{t=1}^T   f_t(\u_t)
\end{equation}
where  $\u_1, \ldots, \u_T \in \X$. Instead of following the definition in (\ref{eqn:dynamic:1}), most of existing studies on dynamic regret consider a restricted form, in which the sequence of comparators consists of local minimizers of online functions \citep{Non-Stationary}, i.e.,
\begin{equation} \label{eqn:dynamic:2}
\begin{split}
R(\x_1^*, \ldots, \x_T^*) =&\sum_{t=1}^T f_t(\x_t)  - \sum_{t=1}^T   f_t(\x_t^*)=\sum_{t=1}^T f_t(\x_t)  -\sum_{t=1}^T \min_{\x \in \X} f_t(\x)
\end{split}
\end{equation}
where $\x_t^* \in \argmin_{\x \in \X} f_t(\x)$ is a minimizer of $f_t(\cdot)$ over domain $\X$. Note that although $R(\u_1, \ldots, \u_T) \leq R(\x_1^*, \ldots, \x_T^*)$, it does not mean the notion of $R(\x_1^*, \ldots, \x_T^*)$ is stronger, because an upper bound for $R(\x_1^*, \ldots, \x_T^*)$ could be very loose for $R(\u_1, \ldots, \u_T)$.

The general dynamic regret in (\ref{eqn:dynamic:1}) includes the static regret in (\ref{eqn:regret}) and the restricted dynamic regret in (\ref{eqn:dynamic:2}) as special cases. Thus, minimizing the general dynamic regret can automatically adapt to the nature of environments, stationary or dynamic. In contrast, the restricted dynamic regret is too pessimistic, and unsuitable for stationary problems. For example, it is meaningless to the problem of statistical machine learning, where  $f_t$'s are sampled independently from the \emph{same} distribution. Due to the random perturbation caused by sampling, the local minimizers could differ significantly from the global minimizer of the expected loss. In this case, minimizing (\ref{eqn:dynamic:2}) will lead to overfitting.

Because of its flexibility, we focus on the general dynamic regret in this paper. Bounding the general dynamic regret is very challenging, because we need to establish a \emph{universal} guarantee that holds for any sequence of comparators. By  comparison,  when bounding the restricted dynamic regret, we only need to focus on the local minimizers. Till now, we  have very limited knowledge on the general dynamic regret. One result is given by \citet{zinkevich-2003-online}, who demonstrates that online gradient descent (OGD) achieves the following dynamic regret bound
\begin{equation} \label{eqn:upper:ogd}
R(\u_1, \ldots, \u_T) = O\left(\sqrt{T} (1+P_T)\right)
\end{equation}
where $P_T$, defined in (\ref{eqn:path}), is the path-length of $\u_1, \ldots, \u_T$.

However, the linear dependence on $P_T$ in (\ref{eqn:upper:ogd}) is too loose, and there is a large gap between the upper bound and the $\Omega(\sqrt{T(1+P_T)})$ lower bound established in our paper. To address this limitation, we propose a novel online method, namely \underline{a}daptive learning for \underline{d}ynamic \underline{e}nvi\underline{r}onment (Ader), which attains an $O(\sqrt{T(1+P_T)})$ dynamic regret. Ader follows the framework of learning with expert advice \citep{bianchi-2006-prediction}, and is inspired by the strategy of maintaining multiple learning rates in MetaGrad \citep{NIPS2016_6268}. The basic idea is to run multiple OGD algorithms in parallel, each with a different step size that is optimal for a specific path-length, and combine them with an expert-tracking algorithm.
While the basic version of Ader needs to query the gradient $O(\log T)$ times in each round, we develop an improved version based on  surrogate loss  and reduce the number of gradient evaluations to $1$. Finally, we provide extensions of Ader to the case that a sequence of dynamical models is given, and obtain tighter bounds when the comparator sequence follows the dynamical models closely.

The contributions of this paper are summarized below.
\begin{compactitem}
\item We establish the \emph{first} lower bound for the general regret bound in (\ref{eqn:dynamic:1}), which is $\Omega(\sqrt{T(1+P_T)})$.
\item We develop a serial of novel methods for minimizing the general dynamic regret, and prove an optimal $O(\sqrt{T(1+P_T)})$ upper bound.
\item Compared to existing work for the restricted dynamic regret in (\ref{eqn:dynamic:2}), our result is \emph{universal} in the sense that the regret bound holds for any sequence of comparators.
\item Our result is also \emph{adaptive} because the upper bound depends on the path-length of the comparator sequence, so it automatically becomes small when  comparators change slowly.
\end{compactitem}
\section{Related Work}
 In this section, we provide a brief review of related work in online convex optimization.

\subsection{Static Regret}
In static setting, online gradient descent (OGD) achieves an $O(\sqrt{T})$ regret bound for general convex functions. If the online functions have additional curvature properties, then faster rates are attainable. For strongly convex functions, the regret bound of OGD becomes $O(\log T)$ \citep{ICML_Pegasos}.  The $O(\sqrt{T})$ and $O(\log T)$ regret bounds, for convex and strongly convex  functions respectively, are known to be minimax optimal \citep{Minimax:Online}. For exponentially concave functions, Online Newton Step (ONS) enjoys an $O(d \log T)$ regret, where $d$ is the dimensionality \citep{ML:Hazan:2007}. When the online functions are both smooth and convex, the regret bound could also be improved if the cumulative loss of the optimal prediction is small \citep{NIPS2010_Smooth}.
\subsection{Dynamic Regret}
To the best of our knowledge, there are only two studies that investigate the general dynamic regret  \citep{zinkevich-2003-online,Dynamic:ICML:13}. While it is impossible to achieve a sublinear dynamic regret in general, we can bound the dynamic regret in terms of certain regularity of the comparator sequence or the function sequence. \citet{zinkevich-2003-online} introduces the path-length
\begin{equation}\label{eqn:path}
P_T(\u_1, \ldots, \u_T)=\sum_{t=2}^T \|\u_t - \u_{t-1}\|_2
\end{equation}
and provides an upper bound for OGD in (\ref{eqn:upper:ogd}).  In a subsequent work, \citet{Dynamic:ICML:13} propose a variant of path-length
\begin{equation}\label{eqn:path:dynamic}
P_T'(\u_1, \ldots, \u_T)=\sum_{t=1}^T \|\u_{t+1} - \Phi_t(\u_{t})\|_2
\end{equation}
in which a sequence of dynamical models $\Phi_t(\cdot): \X \mapsto \X$ is incorporated. Then, they develop a new method, dynamic mirror descent, which achieves an $O(\sqrt{T} (1+P_T'))$ dynamic regret. When the comparator sequence follows the dynamical models closely, $P_T'$ could be much smaller than $P_T$, and thus the upper bound of \citet{Dynamic:ICML:13} could be tighter than that of \citet{zinkevich-2003-online}.

For the restricted dynamic regret, a powerful baseline, which simply plays the minimizer of previous round, i.e., $\x_{t+1} = \argmin_{\x \in \X} f_t(\x)$, attains an $O(P_T^*)$ dynamic regret  \citep{Dynamic:2016}, where
\[
P_T^*:=P_T(\x_1^*, \ldots, \x_T^*)=\sum_{t=2}^T \|\x_t^* - \x_{t-1}^*\|_2.
\]
OGD also achieves the $O(P_T^*)$ dynamic regret, when the online functions are strongly convex and smooth \citep{Dynamic:Strongly}, or when they are convex and smooth and all the minimizers lie in the interior of $\X$ \citep{Dynamic:2016}. Another regularity of the comparator sequence is the squared path-length
\[
S_T^*:=S_T(\x_1^*, \ldots, \x_T^*)=\sum_{t=2}^T \|\x_t^* - \x_{t-1}^*\|_2^2
\]
which could be smaller than the path-length $P_T^*$ when local minimizers move slowly. \citet{Dynamic:Regret:Squared} propose online multiple gradient descent, and establish an $O(\min(P_T^*,S_T^*))$ regret bound for (semi-)strongly convex and smooth functions.

In a recent work, \citet{Non-Stationary} introduce the functional variation
\[
F_T:=F(f_1,\ldots,f_T)= \sum_{t=2}^T \max_{\x \in \X} |f_t(\x) - f_{t-1}(\x)|
\]
to measure the complexity of the function sequence. Under the assumption that an upper bound $V_T \geq F_T$ is known beforehand, \citet{Non-Stationary} develop a  restarted online gradient descent, and prove its dynamic regret is upper bounded by $O(T^{2/3} (V_T+1)^{1/3})$ and $O(\log T \sqrt{T (V_T+1)})$ for convex functions and strongly convex functions, respectively. One limitation of this work is that the bounds are not adaptive because they depend on the upper bound $V_T$. So, even when the actual functional variation $F_T$ is small, the regret bounds do not become better.

One regularity that involves the gradient of functions is
\[
D_T=\sum_{t=1}^T \|\nabla f_t(\x_t)-\m_t\|_2^2
\]
where $\m_1,\ldots,\m_T$ is a predictable sequence  computable by the learner \citep{Gradual:COLT:12,Predictable:COLT:2013}. From the above discussions, we observe that there are different types of regularities. As shown by \citet{Dynamic:AISTATS:15}, these regularities reflect distinct aspects of the online problem, and are not comparable in general. To take advantage of the smaller regularity,  \citet{Dynamic:AISTATS:15} develop an adaptive method whose dynamic regret is on the order of $\sqrt{D_T+1}+\min\{\sqrt{(D_T+1)P_T^*},(D_T+1)^{1/3}T^{1/3}F_T^{1/3}\}$. However, it relies on the assumption that the learner can calculate each regularity online.
\subsection{Adaptive Regret}
Another way to deal with changing environments is to minimize the adaptive regret, which is defined as maximum static regret over any contiguous time interval \citep{Adaptive:Hazan}. For convex functions and exponentially concave functions, \citet{Adaptive:Hazan} have developed efficient algorithms that achieve $O(\sqrt{T \log^3 T})$ and $O(d \log^2 T)$ adaptive regrets, respectively. Later, the adaptive regret of convex functions is improved \citep{Adaptive:ICML:15,Improved:Strongly:Adaptive}. The relation between adaptive regret and restricted dynamic regret is investigated by \citet{Dynamic:Regret:Adaptive}.

\section{Our Methods}
We first state assumptions about the online problem, then provide our motivations, including a lower bound of the general dynamic regret,  and finally present the proposed methods as well as their theoretical guarantees. 
\subsection{Assumptions}
Similar to previous studies in online learning, we introduce the following common assumptions.
\begin{ass}\label{ass:0} On domain $\X$, the values of all functions belong to the range $[a,a+c]$, i.e.,
\[
a \leq f_t(\x) \leq a+c,  \ \forall \x \in \X, \textrm{ and } t \in [T].
\]
\end{ass}
\begin{ass}\label{ass:1} The gradients of all functions are bounded by $G$, i.e.,
\begin{equation}\label{eqn:grad}
\max_{\x \in \X}\|\nabla f_t(\x)\|_2 \leq G, \ \forall t \in [T].
\end{equation}
\end{ass}
\begin{ass}\label{ass:2} The domain $\X$ contains the origin $\mathbf{0}$, and its diameter is bounded by $D$, i.e.,
\begin{equation}\label{eqn:domain}
\max_{\x, \x' \in \X} \|\x -\x'\|_2 \leq D.
\end{equation}
\end{ass}
Note that Assumptions~\ref{ass:1} and \ref{ass:2} imply Assumption \ref{ass:0}  with any $c \geq GD$. In the following, we assume the values of $G$ and $D$ are known to the leaner.

\subsection{Motivations}
According to Theorem 2 of  \citet{zinkevich-2003-online}, we have the following dynamic regret bound for online gradient descent (OGD) with a constant step size.
\begin{thm} \label{thm:ogd}
Consider the online gradient descent (OGD) with $\x_1 \in \X$ and
\[
\x_{t+1} = \Pi_{\X}\big[\x_t - \eta \nabla f_t(\x_t)\big], \ \forall t \geq 1
\]
where $\Pi_{\X}[\cdot]$ denotes the projection onto the nearest point in $\X$.
Under Assumptions~\ref{ass:1} and \ref{ass:2}, OGD satisfies
\[
\begin{split}
\sum_{t=1}^T f_t(\x_t) -  \sum_{t=1}^T f_t(\u_t) \leq   \frac{7 D^2}{4 \eta}  + \frac{D}{\eta} \sum_{t=2}^T\|\u_{t-1}-\u_t\|_2 + \frac{\eta T G^2}{2 }
\end{split}
\]
for \emph{any} comparator sequence $\u_1,\ldots,\u_T \in \X$.
\end{thm}
Thus, by choosing $\eta=O(1/\sqrt{T})$, OGD achieves an $O(\sqrt{T}(1+P_T))$ dynamic regret, that is universal. However, this upper bound is far from the $\Omega(\sqrt{T(1+P_T)})$ lower bound indicated by the theorem below.
\begin{thm} \label{thm:lower} For any online algorithm and any $\tau \in [0, TD]$, there exists a sequence of comparators $\u_1,\ldots,\u_T$ satisfying  Assumption~\ref{ass:2} and a sequence of functions $f_1,\ldots,f_T$ satisfying  Assumption~\ref{ass:1}, such that
\[
 P_T(\u_1, \ldots, \u_T) \leq \tau  \textrm{ and } R(\u_1, \ldots, \u_T)=\Omega\big(G\sqrt{T (D^2 + D \tau)}\big).
\]

\end{thm}
Although there exist lower bounds for the restricted dynamic regret \citep{Non-Stationary,Dynamic:2016}, to the best of our knowledge, this is the \emph{first} lower bound for the general dynamic regret.

Let's drop the universal property for the moment, and suppose we only want to compare against a \emph{specific} sequence $\ub_1,\ldots,\ub_T \in \X$ whose path-length $\Pb_T=\sum_{t=2}^T \|\ub_t - \ub_{t-1}\|_2$ is known beforehand. In this simple setting, we can tune the step size optimally as $\eta^*=O(\sqrt{(1+\Pb_T)/T})$
and obtain an improved $O(\sqrt{T(1+\Pb_T)})$ dynamic regret bound, which matches the lower bound in Theorem~\ref{thm:lower}. 
Thus, when bounding the general dynamic regret, we face the following challenge: On one hand, we want the regret bound to hold for \emph{any} sequence of comparators, but on the other hand, to get a tighter bound, we need to tune the step size for a \emph{specific} path-length. In the next section, we address this dilemma by running multiple OGD algorithms with different step sizes, and combining them through a meta-algorithm.

\subsection{The Basic Approach} \label{sec:basic}
Our proposed method, named as \underline{a}daptive learning for \underline{d}ynamic \underline{e}nvi\underline{r}onment (Ader), is inspired by a recent work for online learning with multiple types of functions---MetaGrad \citep{NIPS2016_6268}.  Ader maintains a set of experts, each attaining an optimal dynamic regret for a different path-length, and chooses the best one using an expert-tracking algorithm.

\paragraph{Meta-algorithm} Tracking the best expert is a well-studied problem \citep{Herbster1998}, and our meta-algorithm, summarized in Algorithm~\ref{alg:1}, is built upon the exponentially weighted average forecaster \citep{bianchi-2006-prediction}. The inputs of the meta-algorithm are its own step size $\alpha$, and a set $\H$ of step sizes for experts. In Step 1, we active a set of experts $\{E^\eta |\eta \in \H\}$ by invoking the expert-algorithm for each $\eta \in \H$. In Step 2, we set the initial weight of each expert. Let $\eta_i$ be the $i$-th smallest step size in $\H$. The weight of $E^{\eta_i}$ is chosen as
\begin{equation}\label{eqn:initial:weight}
w_1^{\eta_i}=\frac{C}{i(i+1)}, \textrm{ and } C=1+\frac{1}{|\H|}.
\end{equation}
In each round, the meta-algorithm receives a set of predictions $\{\x_t^\eta|\eta \in \H\}$ from all experts (Step 4), and  outputs the weighted average (Step 5):
\[
\x_t = \sum_{\eta \in \H} w_t^\eta \x_t^\eta
\]
where $w_t^\eta$ is the weight assigned to expert $E^\eta$. After observing the loss function, the weights of experts are updated according to the  exponential weighting scheme (Step 7):
 \[
 w_{t+1}^\eta = \frac{w_{t}^\eta e^{-\alpha f_t(\x_t^\eta)} }{\sum_{\mu \in \H} w_{t}^\mu e^{-\alpha f_t(\x_t^\mu)}}.
 \]
In the last step, we send the gradient $\nabla f_t(\x_t^\eta)$ to each expert $E^\eta$ so that they can update their own predictions.

\begin{algorithm}[t]
\caption{Ader: Meta-algorithm}
\begin{algorithmic}[1]
\REQUIRE A step size $\alpha$, and a set $\H$ containing  step sizes for experts
\STATE Activate a set of experts $\{E^\eta| \eta \in \H \}$ by invoking Algorithm~\ref{alg:2} for each step size $\eta \in \H$
\STATE Sort step sizes in ascending order $\eta_1 \leq \eta_2 \leq \cdots \leq \eta_N$, and set $w_1^{\eta_i}=\frac{C}{i(i+1)}$
\FOR{$t=1,\ldots,T$}
\STATE Receive $\x_t^\eta$ from each expert $E^\eta$
\STATE Output
\[
\x_t = \sum_{\eta \in \H} w_t^\eta \x_t^\eta
\]
\STATE Observe the loss function $f_t(\cdot)$
\STATE Update the weight of each expert by
\[
w_{t+1}^\eta = \frac{w_{t}^\eta e^{-\alpha f_t(\x_t^\eta)} }{\sum_{\mu \in \H} w_{t}^\mu e^{-\alpha f_t(\x_t^\mu)}}
\]
\STATE Send gradient $\nabla f_t(\x_t^\eta)$ to each expert $E^\eta$
\ENDFOR
\end{algorithmic}
\label{alg:1}
\end{algorithm}
\begin{algorithm}[t]
\caption{Ader: Expert-algorithm}
\begin{algorithmic}[1]
\REQUIRE The step size $\eta$
\STATE Let $\x_1^\eta$ be any point in $\X$
\FOR{$t=1,\ldots,T$}
\STATE Submit $\x_t^\eta$ to the meta-algorithm
\STATE Receive gradient $\nabla f_t(\x_t^\eta)$ from the meta-algorithm
\STATE
\[
\x_{t+1}^\eta= \Pi_{\X}\big[\x_t^\eta - \eta \nabla f_t(\x_t^\eta)\big]
\]
\ENDFOR
\end{algorithmic}
\label{alg:2}
\end{algorithm}

\paragraph{Expert-algorithm} Experts are themselves algorithms, and our expert-algorithm, presented in Algorithm~\ref{alg:2}, is the standard online gradient descent (OGD). Each expert is an instance of OGD, and takes the step size $\eta$ as its input. In Step 3 of Algorithm~\ref{alg:2}, each expert submits its prediction $\x_t^\eta$  to the meta-algorithm, and receives the gradient $\nabla f_t(\x_t^\eta)$ in Step 4. Then, in Step 5 it performs gradient descent
\[
\x_{t+1}^\eta= \Pi_{\X}\big[\x_t^\eta - \eta \nabla f_t(\x_t^\eta)\big]
\]
to get the prediction for the next round.

Next, we specify the parameter setting and our dynamic regret. The set $\H$ is constructed in the way such that for any possible sequence of comparators, there exists a step size that is nearly optimal. To control the size of $\H$, we use a geometric series with ratio $2$. The value of $\alpha$ is tuned such that the upper bound is minimized. Specifically, we have the following theorem.
\begin{thm} \label{thm:ader} Set
\begin{equation} \label{eqn:def:H}
\H=\left\{ \left. \eta_i=\frac{2^{i-1} D}{G}\sqrt{\frac{7}{2T}} \right  | i=1,\ldots, N\right\}
\end{equation}
where $N= \lceil \frac{1}{2} \log_2 (1+4T/7) \rceil +1$, and $\alpha= \sqrt{8 /(T c^2)}$ in Algorithm~\ref{alg:1}.
Under Assumptions~\ref{ass:0}, \ref{ass:1} and \ref{ass:2}, for \emph{any} comparator sequence $\u_1,\ldots,\u_T \in \X$, our proposed Ader method satisfies
\[
\begin{split}
\sum_{t=1}^T f_t(\x_t) -  \sum_{t=1}^T f_t(\u_t) \leq& \frac{3G}{4} \sqrt{2 T (7D^2 + 4 DP_T) }+ \frac{c \sqrt{2T} }{4} \left[1+ 2\ln (k+1)\right] \\
=&O\left(\sqrt{T(1+P_T)} \right)
\end{split}
\]
where
\begin{equation} \label{eqn:def:k}
k = \left\lfloor \frac{1}{2} \log_2 \left(1+ \frac{4P_T}{7D} \right) \right \rfloor+1.
\end{equation}
\end{thm}
The order of the upper bound matches the $\Omega(\sqrt{T(1+P_T)})$ lower bound in Theorem~\ref{thm:lower} exactly.
\subsection{An Improved Approach} \label{sec:improved}
The basic approach in Section~\ref{sec:basic} is simple, but it has an obvious limitation: From Steps 7 and 8 in Algorithm~\ref{alg:1}, we observe that the meta-algorithm needs to query the value and gradient of $f_t(\cdot)$ $N$ times in each round, where $N=O(\log T)$. In contrast, existing algorithms for minimizing static regret, such as OGD, only query the gradient \emph{once} per iteration. When the function is complex, the evaluation of gradients or values could be expensive, and it is appealing to reduce the number of queries in each round.

\paragraph{Surrogate Loss} We introduce \emph{surrogate loss} \citep{NIPS2016_6268} to replace the original loss function. From the first-order condition of convexity  \citep{Convex-Optimization}, we have
\[
f_t(\x) \geq f_t(\x_t)+\langle \nabla f_t(\x_t), \x -\x_t \rangle, \ \forall \x \in \X.
\]
Then, we define the surrogate loss in the $t$-th iteration as
\begin{equation} \label{eqn:surr}
\ell_t(\x)=\langle \nabla f_t(\x_t), \x -\x_t \rangle
\end{equation}
and use it to update the prediction. Because
\begin{equation} \label{eqn:convex}
f_t (\x_t) - f_t (\u_t) \leq \ell_t (\x_t) - \ell_t (\u_t),
\end{equation}
we conclude that the regret w.r.t.~true losses $f_t$'s is smaller than that w.r.t.~surrogate losses $\ell_t$'s. Thus, it is safe to replace $f_t$ with $\ell_t$. The new method, named as improved Ader, is summarized in Algorithms~\ref{alg:3} and \ref{alg:4}.

\paragraph{Meta-algorithm} The new meta-algorithm in Algorithm~\ref{alg:3} differs from the old one in Algorithm~\ref{alg:1} since Step 6. The new algorithm  queries the gradient of $f_t(\cdot)$ at $\x_t$, and then constructs the surrogate loss $\ell_t(\cdot)$ in (\ref{eqn:surr}), which is used in subsequent steps. In Step 8, the weights of experts are updated based on $\ell_t(\cdot)$, i.e.,
 \[
 w_{t+1}^\eta = \frac{w_{t}^\eta e^{-\alpha \ell_t(\x_t^\eta)} }{\sum_{\mu \in \H} w_{t}^\mu e^{-\alpha \ell_t(\x_t^\mu)}}.
 \]
 In Step 9, the gradient of $\ell_t(\cdot)$ at $\x_t^\eta$ is sent to each expert $E^\eta$. Because the surrogate loss is linear,
 \[
 \nabla \ell_t(\x_t^\eta) =  \nabla f_t(\x_t), \ \forall \eta \in \H.
 \]
 As a result, we only need to send the same $\nabla f_t(\x_t)$ to all experts. From the above descriptions, it is clear that the new algorithm only queries the gradient once in each iteration.

\paragraph{Expert-algorithm} The new expert-algorithm in Algorithm~\ref{alg:4} is almost the same as the previous one in Algorithm~\ref{alg:2}. The only difference is that in Step 4, the expert receives the gradient $\nabla f_t(\x_t)$, and uses it to perform gradient descent
\[
\x_{t+1}^\eta= \Pi_{\X}\big[\x_t^\eta - \eta \nabla f_t(\x_t)\big]
\]
in Step 5.

\begin{algorithm}[t]
\caption{Improved Ader: Meta-algorithm}
\begin{algorithmic}[1]
\REQUIRE A step size $\alpha$, and a set $\H$ containing  step sizes for experts
\STATE Activate a set of experts $\{E^\eta| \eta \in \H \}$ by invoking Algorithm~\ref{alg:4} for each step size $\eta \in \H$
\STATE Sort step sizes in ascending order $\eta_1 \leq \eta_2 \leq \cdots \leq \eta_N$, and set $w_1^{\eta_i}=\frac{C}{i(i+1)}$
\FOR{$t=1,\ldots,T$}
\STATE Receive $\x_t^\eta$ from each expert $E^\eta$
\STATE Output
\[
\x_t = \sum_{\eta \in \H} w_t^\eta \x_t^\eta
\]
\STATE Query the gradient of $f_t(\cdot)$ at $\x_t$
\STATE Construct the surrogate loss $\ell_t(\cdot)$ in (\ref{eqn:surr})
\STATE Update the weight of each expert by
\[
w_{t+1}^\eta = \frac{w_{t}^\eta e^{-\alpha \ell_t(\x_t^\eta)} }{\sum_{\mu \in \H} w_{t}^\mu e^{-\alpha \ell_t(\x_t^\mu)}}
\]
\STATE Send gradient $\nabla f_t(\x_t)$ to each expert $E^\eta$
\ENDFOR
\end{algorithmic}
\label{alg:3}
\end{algorithm}
\begin{algorithm}[t]
\caption{Improved Ader: Expert-algorithm}
\begin{algorithmic}[1]
\REQUIRE The step size $\eta$
\STATE Let $\x_1^\eta$ be any point in $\X$
\FOR{$t=1,\ldots,T$}
\STATE Submit $\x_t^\eta$ to the meta-algorithm
\STATE Receive gradient $\nabla f_t(\x_t)$ from the meta-algorithm
\STATE
\[
\x_{t+1}^\eta= \Pi_{\X}\big[\x_t^\eta - \eta \nabla f_t(\x_t)\big]
\]
\ENDFOR
\end{algorithmic}
\label{alg:4}
\end{algorithm}

We have the following theorem to bound the dynamic regret of the improved Ader.
\begin{thm} \label{thm:iader} Use the construction of $\H$ in (\ref{eqn:def:H}), and set $\alpha= \sqrt{2 /(T G^2 D^2)}$ in Algorithm~\ref{alg:3}. Under Assumptions~\ref{ass:1} and \ref{ass:2}, for \emph{any} comparator sequence $\u_1,\ldots,\u_T \in \X$, our improved Ader method satisfies
\[
\begin{split}
\sum_{t=1}^T f_t(\x_t) - \sum_{t=1}^T & f_t(\u_t) \leq  \frac{3G}{4} \sqrt{2 T (7D^2 + 4 DP_T) } + \frac{GD \sqrt{2T}}{2} \left[1+ 2\ln (k+1)\right]\\
=&O\left(\sqrt{T(1+P_T)} \right)
\end{split}
\]
 where $k$ is defined in (\ref{eqn:def:k}).
\end{thm}

Similar to the basic approach, the improved Ader also achieves an $O(\sqrt{T(1+P_T)})$ dynamic regret, that is universal and adaptive. The main advantage is that the improved Ader only needs to query the gradient of the online function \emph{once} in each iteration.

\subsection{Extensions}
Following \citet{Dynamic:ICML:13}, we consider the case that the learner is given a sequence of dynamical models $\Phi_t(\cdot): \X \mapsto \X$, which can be used to characterize the comparators we are interested in. Similar to \citet{Dynamic:ICML:13}, we assume each $\Phi_t(\cdot)$ is a contraction mapping.
\begin{ass}\label{ass:3} All the dynamical models are  contraction mappings, i.e.,
\begin{equation}\label{eqn:phi:con}
\|\Phi_t(\x)-\Phi_t(\x')\|_2 \leq \|\x-\x'\|_2 ,
\end{equation}
for all $t \in [T]$, and $\x, \x' \in \X$.
\end{ass}
Then, we choose $P_T'$ in (\ref{eqn:path:dynamic}) as the regularity of a comparator sequence, which measures how much it deviates from the given dynamics.

\paragraph{Algorithms} For brevity, we only discuss how to incorporate the dynamical models into the basic Ader  in Section~\ref{sec:basic}, and the extension to the improved version can be done in the same way. In fact, we only need to modify the expert-algorithm, and the updated one is provided in Algorithm~\ref{alg:5}. To utilize the dynamical model, after performing gradient descent, i.e.,
\[
\xb_{t+1}^\eta= \Pi_{\X}\big[\x_t^\eta - \eta \nabla f_t(\x_t^\eta)\big]
\]
in Step 5, we apply the dynamical model to the intermediate solution $\xb_{t+1}^\eta$, i.e.,
\[
\x_{t+1}^\eta=\Phi_t(\xb_{t+1}^\eta),
\]
and obtain the prediction for the next round. In the meta-algorithm (Algorithm~\ref{alg:1}), we only need to replace Algorithm~\ref{alg:2} in Step 1 with Algorithm~\ref{alg:5}, and the rest is the same. The dynamic regret of the new algorithm is given below.

\begin{algorithm}[t]
\caption{Ader: Expert-algorithm with dynamical models}
\begin{algorithmic}[1]
\REQUIRE The step size $\eta$, a sequence of dynamical models $\Phi_t(\cdot)$
\STATE Let $\x_1^\eta$ be any point in $\X$
\FOR{$t=1,\ldots,T$}
\STATE Submit $\x_t^\eta$ to the meta-algorithm
\STATE Receive gradient $\nabla f_t(\x_t^\eta)$ from the meta-algorithm
\STATE
\[
\xb_{t+1}^\eta= \Pi_{\X}\big[\x_t^\eta - \eta \nabla f_t(\x_t^\eta)\big]
\]
\STATE
\[
\x_{t+1}^\eta=\Phi_t(\xb_{t+1}^\eta)
\]
\ENDFOR
\end{algorithmic}
\label{alg:5}
\end{algorithm}

\begin{thm} \label{thm:ader:extend} Set
\begin{equation} \label{eqn:def:H:new}
\H=\left\{ \left. \eta_i=\frac{2^{i-1} D}{G}\sqrt{\frac{1}{T}} \right  | i=1,\ldots, N\right\}
\end{equation}
where $N= \left\lceil \frac{1}{2}\log_2 (1+2T)\right\rceil +1$,  $\alpha= \sqrt{8 /(T c^2)}$, and use Algorithm~\ref{alg:5} as the expert-algorithm in Algorithm~\ref{alg:1}. Under Assumptions~\ref{ass:0}, \ref{ass:1}, \ref{ass:2} and \ref{ass:3}, for \emph{any} comparator sequence $\u_1,\ldots,\u_T \in \X$, our proposed Ader method satisfies
\[
\begin{split}
\sum_{t=1}^T f_t(\x_t) - & \sum_{t=1}^T f_t(\u_t) \leq \frac{3G}{2} \sqrt{T (D^2 + 2 DP_T') }+ \frac{c \sqrt{2T} }{4} \left[1+ 2\ln (k+1)\right]\\
=&O\left(\sqrt{T(1+P_T')} \right)
\end{split}
\]
 where
\[
k = \left\lfloor \frac{1}{2} \log_2 \left(1+ \frac{2P_T'}{D}\right) \right\rfloor+1.
\]

\end{thm}
Theorem~\ref{thm:ader:extend} indicates our method achieves an $O(\sqrt{T(1+P_T')})$ dynamic regret, improving the $O(\sqrt{T} (1+P_T'))$ dynamic regret of \citet{Dynamic:ICML:13} significantly. Note that when $\Phi_t(\cdot)$ is the identity map, we recover the result in Theorem~\ref{thm:ader}. Thus, the upper bound in Theorem~\ref{thm:ader:extend} is also  optimal.

\section{Analysis}
In this section, we provide proofs of all the theorems.
\subsection{Proof of Theorem~\ref{thm:ogd}}
For the sake of completeness, we provide the proof here. Let $\x_{t+1}'=\x_t - \eta \nabla f_t(\x_t)$. Following the standard analysis, we have
\begin{equation} \label{eqn:ogd:1}
\begin{split}
 f_t(\x_{t}) - f_t(\u_t)\leq & \langle \nabla f_t(\x_{t}), \x_{t} - \u_t\rangle = \frac{1}{\eta} \langle \x_{t}  - \x_{t+1}', \x_{t} - \u_t\rangle \\
= & \frac{1}{2 \eta} \left( \|\x_t-\u_t\|_2^2 - \|\x_{t+1}'-\u_t\|_2^2 + \|\x_t  - \x_{t+1}'\|_2^2 \right) \\
=& \frac{1}{2 \eta} \left( \|\x_{t}-\u_t\|_2^2 - \|\x_{t+1}'-\u_t\|_2^2 \right) + \frac{\eta}{2 } \|\nabla f_t(\x_{t})\|_2^2  \\
\overset{\text{(\ref{eqn:grad})}}{\leq} & \frac{1}{2 \eta} \left( \|\x_{t}-\u_t\|_2^2 - \|\x_{t+1}-\u_t\|_2^2 \right) + \frac{\eta}{2 } G^2\\
=& \frac{1}{2 \eta} (\|\x_{t}\|_2^2 - \|\x_{t+1}\|_2^2) + \frac{1}{\eta}(\x_{t+1}-\x_t)^\top \u_t + \frac{\eta }{2 } G^2.
\end{split}
\end{equation}
Summing the above inequality over all iterations, we have
\[
\begin{split}
\sum_{t=1}^T \big(f_t(\x_{t}) - f_t(\u_t)\big) \leq & \frac{1}{2 \eta}\|\x_1\|_2^2 + \frac{1}{\eta} \sum_{t=1}^T (\x_{t+1}-\x_t)^\top \u_t + \frac{\eta T}{2 } G^2\\
=& \frac{1}{2 \eta} \|\x_1\|_2^2  + \frac{1}{\eta}(\x_{T+1}^\top \u_T- \x_1^\top \u_1) + \frac{1}{\eta}\sum_{t=2}^T (\u_{t-1}-\u_t)^\top \x_t+ \frac{\eta T}{2 }  G^2 \\
\leq & \frac{7 D^2}{4 \eta}  + \frac{D}{\eta} \sum_{t=2}^T\|\u_{t-1}-\u_t\|_2 + \frac{\eta T}{2 }  G^2
\end{split}
\]
where the last step makes use of
\[
\begin{split}
\|\x_1\|_2^2= \|\x_1-\mathbf{0}\|_2^2 \overset{\text{(\ref{eqn:domain})}}{\leq} D^2, \\
\x_{T+1}^\top \u_T \leq \|\x_{T+1}\|_2\|\u_T\|_2\overset{\text{(\ref{eqn:domain})}}{\leq} D^2,\\
- \x_1^\top \u_1 \leq \frac{1}{4}\|\x_1 -\u_1 \|^2 \overset{\text{(\ref{eqn:domain})}}{\leq} \frac{1}{4} D^2,\\
(\u_{t-1}-\u_t)^\top \x_t \leq \|\u_{t-1}-\u_t\|_2 \|\x_t\|_2 \overset{\text{(\ref{eqn:domain})}}{\leq} D \|\u_{t-1}-\u_t\|_2.
\end{split}
\]
\subsection{Proof of Theorem~\ref{thm:lower}}
Following  \cite{Minimax:Online}, we model online convex optimization as a game between a learner and an adversary, and analyze the minimax value of the dynamic regret.

Let $\X=\{\x:\|\x\|_2 \leq D/2\}$ be a ball with radius $D/2$ which satisfies Assumption~\ref{ass:2}, and $\F$ be the set of convex functions that satisfies Assumption~\ref{ass:1}. We first recall the minimax static regret from \cite{Minimax:Online}:
\[
V_T(X,\F)=\inf_{\x_1 \in \X} \sup_{f_1 \in \F} \cdots \inf_{\x_T \in \X} \sup_{f_T \in \F} \left( \sum_{t=1}^T f_t(\x_t) - \min_{\x \in \X} \sum_{t=1}^T f_t(\x)\right)=\frac{GD}{2}\sqrt{T}.
\]
Given any path-length $\tau \in [0, TD]$, we define the set of comparator sequences whose path-lengths are no more than $\tau$ as
\[
\C(\tau)=\left\{\u_1,\ldots,\u_T \in \X: \sum_{t=2}^T \|\u_t - \u_{t-1}\|_2 \leq \tau \right\}.
\]
Then, the minimax dynamic regret w.r.t.~$\C(\tau)$ is defined as
\[
V_T(X,\F,\C(\tau))=\inf_{\x_1 \in \X} \sup_{f_1 \in \F} \cdots \inf_{\x_T \in \X} \sup_{f_T \in \F} \left( \sum_{t=1}^T f_t(\x_t) - \min_{\u_1,\ldots,\u_T \in \C(\tau)} \sum_{t=1}^T f_t(\u_t)\right).
\]

We consider two cases: $\tau < D$ and $\tau \geq D$.  When $\tau < D$, we use the minimax static regret to lower bound the minimax dynamic regret:
\begin{equation}\label{eqn:lower:1}
V_T(X,\F,\C(\tau)) \geq V_T(X,\F)=\frac{GD}{2}\sqrt{T}.
\end{equation}

Next, we consider the case $\tau \geq D$. Without loss of generality, we assume $\lceil \tau/D \rceil$ divides $T$, and define $L= T/\lceil \tau/D \rceil$. To proceed, we construct $\C'(\tau)$, a subset of $\C(\tau)$, which keeps the comparator fixed for each successive $L$ rounds, that is,
\[
\C'(\tau)=\Big\{\u_1,\ldots,\u_T \in \X: \u_{(i-1)L+1}=\u_{(i-1)L+2} =\cdots= \u_{iL} , \forall i \in \big[1,  \lceil \tau/D \rceil\big] \Big\}.
\]
Since each sequence in $\C'(\tau)$ changes at most $\lceil \tau/D \rceil-1 \leq \tau/D$ times, its path-length does not exceed $\tau$. As a result,  $\C'(\tau)  \subseteq \C(\tau)$ implying
\begin{equation} \label{eqn:lower:2}
V_T(X,\F,\C(\tau)) \geq V_T(X,\F,\C'(\tau)).
\end{equation}
The advantage of introducing $\C'(\tau)$ is that the minimax dynamic regret w.r.t.~$\C'(\tau)$ can be decomposed as the sum of $\lceil \tau/D \rceil$ minimax static regret of length $L$. Specifically, we have
\begin{equation}\label{eqn:lower:3}
\begin{split}
 V_T(X,\F,\C'(\tau))= &\inf_{\x_1 \in \X} \sup_{f_1 \in \F} \cdots \inf_{\x_T \in \X} \sup_{f_T \in \F} \left( \sum_{t=1}^T f_t(\x_t) - \min_{\u_1,\ldots,\u_T \in \C'(\tau)} \sum_{t=1}^T f_t(\u_t)\right)\\
=& \inf_{\x_1 \in \X} \sup_{f_1 \in \F} \cdots \inf_{\x_T \in \X} \sup_{f_T \in \F} \left( \sum_{t=1}^T f_t(\x_t) - \sum_{i=1}^{\lceil \tau/D \rceil} \min_{\x \in \X} \sum_{t=(i-1)L+1}^{iL} f_t(\x) \right)\\
=& \lceil \tau/D \rceil  \frac{GD}{2}\sqrt{L} =  \frac{GD}{2}\sqrt{T \lceil  \tau/D \rceil } \geq \frac{G}{2} \sqrt{T D \tau}.
\end{split}
\end{equation}
Substituting  (\ref{eqn:lower:3}) into (\ref{eqn:lower:2}), we have
\begin{equation} \label{eqn:lower:4}
V_T(X,\F,\C(\tau)) \geq  \frac{G}{2} \sqrt{T D \tau}.
\end{equation}

Combining (\ref{eqn:lower:1}) and (\ref{eqn:lower:4}), we have
\[
V_T(X,\F,\C(\tau)) \geq \frac{G}{2} \max\left(D \sqrt{T},  \sqrt{T D \tau} \right) =\Omega\big(G\sqrt{T (D^2 + D \tau)}\big)
\]
which completes the proof.
\subsection{Proof of Theorem~\ref{thm:ader}}
The analysis is divided into three parts. First, we show that the cumulative loss of the meta-algorithm is comparable to all  experts. Then, we demonstrate that for any sequence of comparators, there is an expert whose dynamic regret is almost optimal. Finally, putting the regret bounds of the meta-algorithm and experts together, we obtain the dynamic regret of our Ader method.

Following the analysis of exponentially weighted average forecaster, we bound the regret of the meta-algorithm with respect to all experts simultaneously.
\begin{lemma} \label{lem:meta} Under Assumption~\ref{ass:0}, the  regret of  Algorithm~\ref{alg:1} satisfies
\[
\sum_{t=1}^T f_t(\x_t) - \min_{\eta \in \H}\left( \sum_{t=1}^T f_t(\x_t^\eta)  + \frac{1}{\alpha} \ln \frac{1}{w_1^\eta} \right)  \leq \frac{\alpha T c^2}{8}.
\]
\end{lemma}
Then, by choosing $\alpha=\sqrt{8/(T c^2})$ to minimize the upper bound, we have
\begin{equation}\label{thm:main:1}
\sum_{t=1}^T f_t(\x_t)  - \sum_{t=1}^T f_t(\x_t^\eta) \leq \frac{c \sqrt{2T}}{4}\left(1+ \ln \frac{1}{w_1^\eta}\right)
\end{equation}
for any $\eta \in \H$.

Next, consider a sequence of comparators $\u_1,\ldots,\u_T \in \X$. Recall that each expert $E^\eta$ performs online gradient descent with step size $\eta$. From Theorem~\ref{thm:ogd}, for each $\eta \in \H$, we have
\begin{equation}\label{thm:main:2}
\sum_{t=1}^T f_t(\x_t^\eta) -  \sum_{t=1}^T f_t(\u_t) \leq  \frac{7 D^2}{4 \eta}  + \frac{D P_T}{\eta}  + \frac{\eta T G^2}{2}.
\end{equation}
Then, we just need to show there is an $\eta_k \in \H$ such that the R.H.S.~of (\ref{thm:main:2}) is almost minimal. If we minimize the R.H.S.~of (\ref{thm:main:2}) exactly, the optimal step size is
\begin{equation}\label{thm:main:3}
\eta^*(P_T)= \sqrt{\frac{7 D^2+ 4 D P_T}{2T G^2} }.
\end{equation}
From Assumption~\ref{ass:2}, we have the following bound of the path-length
\[
0 \leq P_T=\sum_{t=2}^T \|\u_t - \u_{t-1}\|_2 \overset{\text{(\ref{eqn:domain})}}{\leq} TD.
\]
Thus
\[
\frac{D}{G} \sqrt{\frac{7}{2T} } \leq \eta^*(P_T) \leq \frac{D}{G} \sqrt{\frac{7}{2T}+2}.
\]
From our construction of $\H$ in (\ref{eqn:def:H}), it is easy to verify that
\[
\min \H= \frac{D}{G} \sqrt{\frac{7}{2T} } , \textrm{ and  } \max \H \geq \frac{D}{G} \sqrt{\frac{7}{2T}+2}.
\]
As a result, for any possible value of $P_T$, there exists a step size $\eta_k \in \H$, such that
\begin{equation}\label{thm:main:4}
\eta_k =\frac{2^{k-1} D}{G}\sqrt{\frac{7}{2T}} \leq \eta^*(P_T) \leq 2 \eta_k
\end{equation}
where $k = \lfloor \frac{1}{2} \log_2 (1+ \frac{4P_T}{7D}) \rfloor+1$.

Plugging $\eta_k$ into (\ref{thm:main:2}), the regret bound of expert $E^{\eta_k}$ is given by
\begin{equation}\label{thm:main:5}
\begin{split}
&\sum_{t=1}^T f_t(\x_t^{\eta_k}) -  \sum_{t=1}^T f_t(\u_t) \\
\leq & \frac{7 D^2}{4 \eta_k}  + \frac{D P_T}{\eta_k}  + \frac{\eta_k T G^2}{2}\\
\overset{\text{(\ref{thm:main:4})}}{\leq} & \frac{7 D^2}{2 \eta^*(P_T)}  + \frac{2 D P_T}{\eta^*(P_T)}  + \frac{\eta^*(P_T) T G^2}{2}\\
\overset{\text{(\ref{thm:main:3})}}{=}& \frac{3G}{4} \sqrt{2 T (7D^2 + 4 DP_T)}.
\end{split}
\end{equation}

From (\ref{eqn:initial:weight}), we know the initial weight of expert $E^{\eta_k}$ is
\[
w_1^{\eta_k}=\frac{C}{k(k+1)} \geq \frac{1}{k(k+1)} \geq \frac{1}{(k+1)^2}.
\]
Combining with (\ref{thm:main:1}), we obtain the regret of the meta-algorithm w.r.t.~expert $E^{\eta_k}$
\begin{equation}\label{thm:main:6}
\sum_{t=1}^T f_t(\x_t)  - \sum_{t=1}^T f_t(\x_t^{\eta_k}) \leq \frac{c \sqrt{2T} }{4} \left[1+ 2\ln (k+1)\right].
\end{equation}

Finally, from  (\ref{thm:main:5}) (\ref{thm:main:6}), we derive the following dynamic regret bound of Ader
\[
\begin{split}
\sum_{t=1}^T f_t(\x_t) -  \sum_{t=1}^T f_t(\u_t) \leq \frac{3G}{4} \sqrt{2 T (7D^2 + 4 DP_T) }+ \frac{c \sqrt{2T} }{4} \left[1+ 2\ln (k+1)\right]
\end{split}
\]
which holds for any sequence of comparators.

\subsection{Proof of Lemma~\ref{lem:meta}}
Following pervious studies \citep[Theorem 2.2 and Exercise 2.5]{bianchi-2006-prediction}, we define
\[
L_t^\eta= \sum_{i=1}^t f_i(\x_i^\eta), \textrm{ and } W_t= \sum_{\eta \in \H} w_1^\eta e^{- \alpha L_t^\eta}.
\]
From the updating rule in Algorithm~\ref{alg:1}, it is easy to verify that
\begin{equation} \label{eqn:wt:eta}
w_t^\eta =  \frac{w_1^\eta  e^{-\alpha L_{t-1}^\eta}}{\sum_{\mu \in \H} w_1^\mu  e^{-\alpha L_{t-1}^\mu} } , \ t \geq 2.
\end{equation}

First, we have
\begin{equation} \label{eqn:wt:w0}
\begin{split}
\ln W_T = \ln \left( \sum_{\eta \in \H} w_1^\eta e^{- \alpha L_T^\eta}\right) \geq \ln\left( \max_{\eta \in \H}  w_1^\eta e^{- \alpha L_T^\eta} \right) =- \alpha \min_{\eta \in \H} \left(L_T^\eta + \frac{1}{\alpha} \ln \frac{1}{w_1^\eta} \right) .
\end{split}
\end{equation}

Next, we bound the related quantity $\ln(W_t/W_{t-1})$ as follows. When $t \geq 2$, we have
\begin{equation} \label{eqn:wt:wt1}
\begin{split}
&\ln \left(\frac{W_t}{W_{t-1}} \right)= \ln \left( \frac{\sum_{\eta \in \H} w_1^\eta e^{- \alpha L_t^\eta}}{\sum_{\eta \in \H} w_1^\eta e^{- \alpha L_{t-1}^\eta}} \right) \\
=& \ln \left( \frac{\sum_{\eta \in \H} w_1^\eta e^{- \alpha L_{t-1}^\eta} e^{-\alpha f_t(\x_t^\eta) }}{\sum_{\eta \in \H} w_1^\eta e^{- \alpha L_{t-1}^\eta}} \right) \overset{\text{(\ref{eqn:wt:eta})}}{=}  \ln \left(  \sum_{\eta \in \H} w_t^\eta  e^{-\alpha f_t(\x_t^\eta) } \right).
\end{split}
\end{equation}
When $t=1$, we have
\begin{equation} \label{eqn:w1}
\ln W_1=  \ln \left(  \sum_{\eta \in \H} w_1^\eta  e^{-\alpha f_1(\x_1^\eta) } \right).
\end{equation}
Thus
\begin{equation} \label{eqn:WT}
\ln W_T= \ln W_1 +\sum_{t=2}^T\ln \left(\frac{W_t}{W_{t-1}}\right) \overset{\text{(\ref{eqn:wt:wt1}),(\ref{eqn:w1})}}{=}\sum_{t=1}^T \ln \left(  \sum_{\eta \in \H} w_t^\eta  e^{-\alpha f_t(\x_t^\eta) } \right).
\end{equation}

To proceed, we introduce Hoeffding's inequality \citep{Hoeffding}.
\begin{lemma} \label{lem:hoeffding} Let $X$ be a random variable with $a\leq X \leq b$. Then, for any $s \in \R$,
\[
\ln \E\left[e^{sX} \right]\leq s \E[X]+\frac{s^2(b-a)^2}{8}.
\]
\end{lemma}
From Lemma~\ref{lem:hoeffding} and Assumption~\ref{ass:0}, we have
\begin{equation} \label{eqn:hoeffding}
\begin{split}
\ln \left(  \sum_{\eta \in \H} w_t^\eta  e^{-\alpha f_t(\x_t^\eta) } \right) \leq & -\alpha  \sum_{\eta \in \H} w_t^\eta  f_t(\x_t^\eta)+\frac{\alpha^2c^2}{8} \\
\leq & -\alpha  f_t \left( \sum_{\eta \in \H}  w_t^\eta \x_t^\eta \right)+\frac{\alpha^2c^2}{8}
=  -\alpha  f_t \left( \x_t\right)+\frac{\alpha^2c^2}{8}
\end{split}
\end{equation}
where the second inequality is due to Jensen's inequality \citep{Convex-Optimization}.

Substituting (\ref{eqn:hoeffding}) into (\ref{eqn:WT}), we obtain
\[
\ln W_T \leq -\alpha  \sum_{t=1}^T f_t \left( \x_t\right)+\frac{ T \alpha^2c^2}{8}.
\]
Combining this with the lower bound in (\ref{eqn:wt:w0}), we have
\[
- \alpha \min_{\eta \in \H} \left(L_T^\eta + \frac{1}{\alpha} \ln \frac{1}{w_1^\eta} \right) \leq -\alpha  \sum_{t=1}^T f_t \left( \x_t\right)+\frac{ T \alpha^2c^2}{8}.
\]
We complete the proof by simplifying the above inequality.

\subsection{Proof of Theorem~\ref{thm:iader}}
The proof is similar to that of Theorem~\ref{thm:ader}. We just need to replace the original function $f_t(\cdot)$ with the surrogate loss $\ell_t(\cdot)$, and obtain a dynamic regret in terms of the surrogate losses. Then, based on the relation between $f_t(\cdot)$ and $\ell_t(\cdot)$ in (\ref{eqn:convex}), we obtain the dynamic regret in terms of the original functions.

First, we bound the regret of the meta-algorithm with respect to all experts simultaneously. To get a counterpart of Lemma~\ref{lem:meta}, we need to bound the value of $\ell_t(\cdot)$. From Assumptions~\ref{ass:1} and \ref{ass:2}, we have
\[
\begin{split}
|\ell_t(\x)|=|\langle \nabla f_t(\x_t), \x -\x_t \rangle| \leq  \|\nabla f_t(\x_t)\|_2 \|\x -\x_t\|_2 \overset{\text{(\ref{eqn:grad}),(\ref{eqn:domain})}}{\leq}  GD, \ \forall \x \in \X,  t \in [T].
\end{split}
\]
Then, following the same analysis as Lemma~\ref{lem:meta}, we have the following result.
\begin{lemma} \label{lem:meta:new}  Under Assumptions~\ref{ass:1} and \ref{ass:2}, the regret of  Algorithm~\ref{alg:3} satisfies
\[
\sum_{t=1}^T \ell_t(\x_t) - \min_{\eta \in \H}\left( \sum_{t=1}^T \ell_t(\x_t^\eta)  + \frac{1}{\alpha} \ln \frac{1}{w_1^\eta} \right) \leq \frac{\alpha T G^2 D^2}{2}.
\]
\end{lemma}
Then, by choosing $\alpha=\sqrt{2 /(T G^2D^2})$ to minimize the upper bound, we have
\begin{equation}\label{thm:new:1}
\sum_{t=1}^T \ell_t(\x_t)  - \sum_{t=1}^T \ell_t(\x_t^\eta) \leq \frac{GD \sqrt{2T}}{2} \left(1+ \ln \frac{1}{w_1^\eta}\right)
\end{equation}
for any $\eta \in \H$.

Next, since $\ell_t(\cdot)$ is convex, and its gradient is bounded by $G$, the derivation of (\ref{thm:main:5}) can be followed with $f_t(\cdot)$ replaced by $\ell_t(\cdot)$. Specifically, we have
\begin{equation}\label{thm:new:4}
\begin{split}
\sum_{t=1}^T \ell_t(\x_t^{\eta_k}) -  \sum_{t=1}^T \ell_t(\u_t) \leq \frac{3G}{4} \sqrt{2 T (7D^2 + 4 DP_T) }.
\end{split}
\end{equation}
for certain $\eta_k \in \H$ and $w_1^{\eta_k} \geq \frac{1}{(k+1)^2}$.

From (\ref{thm:new:1}), we get the regret of the meta-algorithm w.r.t.~expert $E^{\eta_k}$
\begin{equation}\label{thm:new:5}
\sum_{t=1}^T \ell_t(\x_t)  - \sum_{t=1}^T \ell_t(\x_t^{\eta_k}) \leq \frac{GD \sqrt{2T}}{2} \left[1+ 2\ln (k+1)\right].
\end{equation}

Combining (\ref{thm:new:4}) with (\ref{thm:new:5}), we have
\[
\begin{split}
\sum_{t=1}^T \ell_t(\x_t)-\sum_{t=1}^T \ell_t(\u_t)\leq  \frac{3G}{4} \sqrt{2 T (7D^2 + 4 DP_T) } + \frac{GD \sqrt{2T}}{2} \left[1+ 2\ln (k+1)\right]
\end{split}
\]
which holds for any sequence of comparators. We complete the proof by noticing
\[
\begin{split}
\sum_{t=1}^T f_t(\x_t) -  \sum_{t=1}^T f_t(\u_t) \overset{\text{(\ref{eqn:convex})}}{\leq } \sum_{t=1}^T \ell_t(\x_t) -  \sum_{t=1}^T \ell_t(\u_t).
\end{split}
\]
\subsection{Proof of Theorem~\ref{thm:ader:extend}}
Again, the proof is similar to that of Theorem~\ref{thm:ader}.  Since the meta-algorithm is the same, its regret bound in (\ref{thm:main:6}) can be reused directly. We only need to update the theoretical guarantees of experts. To this end, we develop the following theorem, which is a counterpart of Theorem~\ref{thm:ogd} when dynamical models are incorporated.
\begin{thm} \label{thm:ogd:new}
Consider the following updating rules which incorporate dynamical models into online gradient descent:
\[
\begin{split}
\xb_{t+1} = &\Pi_{\X}\big[\x_t - \eta \nabla f_t(\x_t)\big],\\
\x_{t+1}=&\Phi_t(\xb_{t+1}), \ \forall t \geq 1
\end{split}
\]
where $\x_1 \in \X$ and $\Pi_{\X}[\cdot]$ denotes the projection onto the nearest point in $\X$.
Under Assumptions~\ref{ass:1}, \ref{ass:2} and \ref{ass:3}, it satisfies
\[
\begin{split}
\sum_{t=1}^T f_t(\x_t) -  \sum_{t=1}^T f_t(\u_t) \leq  \frac{D^2}{2 \eta}  + \frac{D}{\eta} \sum_{t=1}^T \|\u_{t+1}-\Phi_t(\u_t)\|_2 + \frac{\eta T G^2}{2 }
\end{split}
\]
for \emph{any} comparator sequence $\u_1,\ldots,\u_T \in \X$.
\end{thm}

Then, we repeat the arguments in Theorem~\ref{thm:ader} to bound the regret of experts. From Theorem~\ref{thm:ogd:new}, for each $\eta \in \H$, we have
\begin{equation}\label{thm:main:2:new}
\sum_{t=1}^T f_t(\x_t^\eta) -  \sum_{t=1}^T f_t(\u_t) \leq  \frac{D^2}{2\eta}  + \frac{D P_T'}{\eta}  + \frac{\eta T G^2}{2}.
\end{equation}
Next, we identify an $\eta_k$ from $\H$ such that the R.H.S.~of (\ref{thm:main:2:new}) is almost minimal. If we minimize the R.H.S.~of (\ref{thm:main:2:new}) exactly, the optimal step size is
\begin{equation}\label{thm:main:3:new}
\eta^*(P_T')= \sqrt{\frac{D^2+ 2 D P_T'}{T G^2} }.
\end{equation}
From Assumption~\ref{ass:2}, we have the following bound of $P_T'$
\[
0 \leq P_T'= \sum_{t=1}^T \|\u_{t+1} - \Phi_t(\u_{t})\|_2\overset{\text{(\ref{eqn:domain})}}{\leq} TD.
\]
Thus
\[
\frac{D}{G} \sqrt{\frac{1}{T} } \leq \eta^*(P_T') \leq \frac{D}{G} \sqrt{\frac{1}{T}+2}.
\]
From our construction of $\H$ in (\ref{eqn:def:H:new}), it is easy to verify that
\[
\min \H= \frac{D}{G} \sqrt{\frac{1}{T} } , \textrm{ and  } \max \H \geq \frac{D}{G} \sqrt{\frac{1}{T}+2}.
\]
As a result, for any possible value of $P_T'$, there must exist a step size $\eta_k \in \H$, such that
\begin{equation}\label{thm:main:4:new}
\eta_k =\frac{2^{k-1} D}{G}\sqrt{\frac{1}{T}}  \leq \eta^*(P_T') \leq 2 \eta_k
\end{equation}
where $k = \lfloor \frac{1}{2} \log_2 (1+ \frac{2P_T'}{D}) \rfloor+1$.

Plugging $\eta_k$ into (\ref{thm:main:2:new}), we have
\begin{equation}\label{thm:main:5:new}
\begin{split}
&\sum_{t=1}^T f_t(\x_t^{\eta_k}) -  \sum_{t=1}^T f_t(\u_t) \leq  \frac{ D^2}{2 \eta_k}  + \frac{D P_T'}{\eta_k}  + \frac{\eta_k T G^2}{2}\\
\overset{\text{(\ref{thm:main:4:new})}}{\leq} & \frac{D^2}{ \eta^*(P_T')}  + \frac{2 D P_T'}{\eta^*(P_T')}  + \frac{\eta^*(P_T) T G^2}{2}\overset{\text{(\ref{thm:main:3:new})}}{=} \frac{3G}{2} \sqrt{T (D^2 + 2 DP_T') }.
\end{split}
\end{equation}
We complete the proof by combining (\ref{thm:main:6}) with (\ref{thm:main:5:new}).

\subsection{Proof of Theorem~\ref{thm:ogd:new}}
Following the derivation of (\ref{eqn:ogd:1}), we have
\[
\begin{split}
 &f_t(\x_{t}) - f_t(\u_t)\\
 \leq &  \frac{1}{2 \eta} \left( \|\x_{t}-\u_t\|_2^2 - \|\xb_{t+1}-\u_t\|_2^2 \right) + \frac{\eta}{2 } G^2\\
\overset{\text{(\ref{eqn:phi:con})}}{\leq} &\frac{1}{2 \eta} \left( \|\x_{t}-\u_t\|_2^2 - \|\Phi_t(\xb_{t+1})-\Phi_t(\u_t)\|_2^2 \right) + \frac{\eta}{2 } G^2\\
= &\frac{1}{2 \eta} \left( \|\x_{t}-\u_t\|_2^2 -\|\x_{t+1}-\u_{t+1}\|_2^2 + \|\x_{t+1}-\u_{t+1}\|_2^2-\|\x_{t+1}-\Phi_t(\u_t)\|_2^2 \right) + \frac{\eta}{2 } G^2\\
=&\frac{1}{2 \eta} \left( \|\x_{t}-\u_t\|_2^2 -\|\x_{t+1}-\u_{t+1}\|_2^2 + \big(\x_{t+1}-\u_{t+1}+\x_{t+1}-\Phi_t(\u_t)\big)^\top (\Phi_t(\u_t)-\u_{t+1}) \right) + \frac{\eta}{2 } G^2\\
\leq & \frac{1}{2 \eta} \left( \|\x_{t}-\u_t\|_2^2 -\|\x_{t+1}-\u_{t+1}\|_2^2 + \big(\|\x_{t+1}-\u_{t+1}\|_2 + \|\x_{t+1}-\Phi_t(\u_t)\|_2 \big)\|\u_{t+1}-\Phi_t(\u_t)\|_2 \right) \\
&+ \frac{\eta}{2 } G^2\\
\overset{\text{(\ref{eqn:domain})}}{\leq} & \frac{1}{2 \eta} \left( \|\x_{t}-\u_t\|_2^2 -\|\x_{t+1}-\u_{t+1}\|_2^2 \right) + \frac{D}{\eta} \|\u_{t+1}-\Phi_t(\u_t)\|_2 + \frac{\eta}{2 } G^2.
\end{split}
\]
Summing the above inequality over all iterations, we have
\[
\begin{split}
\sum_{t=1}^T \big(f_t(\x_{t}) - f_t(\u_t)\big) \leq & \frac{1}{2 \eta}\|\x_1-\u_1\|_2^2 + \frac{D}{\eta} \sum_{t=1}^T \|\u_{t+1}-\Phi_t(\u_t)\|_2 + \frac{\eta T}{2 } G^2\\
\overset{\text{(\ref{eqn:domain})}}{\leq} &  \frac{1}{2 \eta} D^2 + \frac{D}{\eta} \sum_{t=1}^T \|\u_{t+1}-\Phi_t(\u_t)\|_2 + \frac{\eta T}{2 } G^2.
\end{split}
\]

\section{Conclusion and Future Work}
In this paper, we study the general form of dynamic regret, which compares the  cumulative loss of the online learner against an arbitrary sequence of comparators. To this end, we develop a novel method, named as \underline{a}daptive learning for \underline{d}ynamic \underline{e}nvi\underline{r}onment (Ader). Theoretical analysis shows that Ader achieves an optimal $O(\sqrt{T(1+P_T)})$ dynamic regret. When a sequence of dynamical models is available, we extend Ader to incorporate this additional information, and obtain an $O(\sqrt{T(1+P_T')})$ dynamic regret.

In the future, we will investigate whether the curvature of functions, such as strong convexity and smoothness, can be utilized to improve the dynamic regret bound. We note that in the setting of the restricted dynamic regret, the curvature of functions indeed makes the upper bound tighter \citep{Dynamic:Strongly,Dynamic:Regret:Squared}. But whether it improves the general dynamic regret remains an open problem.

\bibliography{E:/MyPaper/ref}
\bibliographystyle{abbrvnat}
\end{document}